A 0.00453 L of sodium chloride solution was needed in a 560 mg sample that assayed


87.1% (w/w) Na2SO4. Calculate the molar concentration of sodium chloride in the


solution.

1
Expert's answer
2022-05-25T09:52:02-0400

C = n(solute) / V(solution)

V(Na2SO4) = m(Na2SO4) / ρ(Na2SO4) = (0.56 g * 87.1 %) / 2.66 g/cm3 = 18.34 cm3 = 0.01834 L

n(NaCl) = m(NaCl) / Mr(NaCl) = (ρ(NaCl)*V(NaCl)) / Mr(NaCl) = (2.16 g/cm3 * 4.53 cm3) / 58,44 g/mol = 0.17 moles


C = 0.17 moles / 0.01834 L = 9.27 mol/L